### v1

# Q-PETR: Quant-aware Position Embedding Transformation for Multi-View 3D Object Detection

Abstract
PETR-based methods have dominated benchmarks in 3D perception and are increasingly becoming a key component in modern autonomous driving systems. However, their quantization performance significantly degrades when INT8 inference is required, with a degradation of 58.2% in mAP and 36.9% in NDS on the NuScenes dataset. To address this issue, we propose a quantization-aware position embedding transformation for multi-view 3D object detection, termed Q-PETR. Q-PETR offers a quantization-friendly and deployment-friendly architecture while preserving the original performance of PETR. It substantially narrows the accuracy gap between INT8 and FP32 inference for PETR-series methods. Without bells and whistles, our approach reduces the mAP and NDS drop to within 1% under standard 8-bit per-tensor post-training quantization. Furthermore, our method exceeds the performance of the original PETR in terms of floating-point precision. Extensive experiments across a variety of PETR-series models demonstrate its broad generalization.
1 Introduction
3D perception is crucial for autonomous driving and embodied intelligence [17, 67, 31, 16, 62]. PETR-based methods [49, 29, 30, 43, 56, 50, 52] have gained significant attention due to their effectiveness in this domain. Unlike previous detection methods relying on dense convolutional features [17, 61, 60, 59, 37] or utilizing 3D-to-2D projections as in DETR3D-based approaches [51, 25, 28], PETR adapts the paradigm of 2D object detection with transformers (DETR) [5] by incorporating 3D positional encoding, achieving seamless end-to-end 3D obstacle detection.
Despite its advantages, PETR requires substantial storage and computational resources during inference, posing challenges for on-device deployment. To enhance inference efficiency, considerable efforts have been dedicated to efficient model compression [42, 20, 27, 19, 11, 12], with quantization being particularly favored for deploying models on AI chips by representing networks in low-bit formats [39, 32, 54, 7].
However, deploying complex models like PETR on edge devices presents additional challenges. Autonomous vehicles rely on edge AI chips with limited computational capabilities. Nonlinear functions such as softmax, sigmoid, SiLU, and GeLU are commonly used during inference, but most edge AI chips lack sufficient floating-point performance for these operations, exacerbating deployment difficulties.
Moreover, quantization awareness is often overlooked in neural architecture design, leading to significant performance degradation during deployment. Notably, models like MobileNet [40, 64, 15, 38], EfficientNet [4, 13, 45], and RepVGG [8] face quantization issues requiring additional design efforts or advanced strategies to mitigate deployment challenges. PETR is no exception; when deployed on in-vehicle AI chips, standard 8-bit per-tensor post-training quantization (PTQ) results in a significant performance drop, with a 58.2% decrease in mAP and a 36.9% decrease in NDS.
In this paper, we address the quantization issues of PETR to facilitate its deployment on edge AI chips in autonomous vehicles. We conduct an in-depth analysis of the mechanisms behind PETR’s quantization failure, revealing that the significantly larger magnitudes of positional encodings compared to image features, and the highly imbalanced scaled dot-products in cross-attention, are the primary causes of quantization collapse. Based on these insights, we propose Q-PETR (Quantization-aware PETR), a model that mitigates quantization collapse and even improves floating-point performance compared to the original PETR.
Our contributions are summarized as follows:
Identifying the root causes of quantization collapse in PETR: We uncover that the imbalanced output of the inverse-sigmoid function and excessively large values in positional encoding, as well as the imbalanced scaled dot-products in cross-attention, are key issues.
Redesigning the positional encoding and quantization strategy: We reformulate the positional encoding and improve the quantization approach for the scaled dot-product in cross-attention, enhancing both floating-point performance and quantization friendliness.
Introducing DuLUT for efficient nonlinear function inference: We propose DuLUT, an improved lookup table [46] approach that reduces the number of table entries without compromising precision, optimizing nonlinear function inference on edge AI chips.
Demonstrating generalizability and deployment readiness: Our method generalizes well across different model scales, achieving exceptional post-quantization performance suitable for deployment on edge devices.
2 Related Work
3D Object Detection. Surround-view 3D object detection is essential for autonomous driving and is generally categorized into LSS-based [17, 59, 16] and transformer-based [29, 43] approaches. LSS-based methods project multi-camera features onto dense BEV (Bird’s Eye View) representations [37], but their high memory consumption hinders efficient long-range perception. Transformer-based methods leverage sparsity to enhance long-distance perception. Among these, the PETR series has gained significant attention. PETR [29] transforms 2D image features into 3D representations using 3D positional encoding. PETRv2 [30] introduces temporal feature indexing, while StreamPETR [49] extends temporal query processing. Some works [48, 52, 9] accelerate processing by incorporating 2D detection priors. CMT [56] fuses vision and LiDAR point clouds. Improvements to PETR’s positional encoding have also been explored [43, 14]. Additionally, PETR has been integrated into the Omnidrive framework [50] to enhance 3D perception with large models.
Quantization. Quantization compresses models by converting weights and activations from floating-point to lower-bit integer representations [6, 65, 3, 7]. Among various methods [39, 53, 32, 1, 2, 54], we focus on uniform symmetric quantization, mapping floating-point values to discrete -bit integer values as:
where is the scaling factor computed as:
with and being the maximum and minimum floating-point values from the calibration dataset. Quantization methods are categorized into Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). QAT [10, 4] introduces quantization-aware losses during training, enhancing robustness but requiring resource-intensive retraining. PTQ offers rapid deployment without retraining. While PTQ methods have been successful on CNNs [33, 34, 21], they often perform poorly on Vision Transformers (ViTs) and transformer-based detectors due to structural differences. For ViTs, practical PTQ algorithms have been developed [63, 26, 44, 24, 41]. In the context of transformer-based object detection models, Q-DETR [55] and AQ-DETR [47] use QAT and knowledge distillation to mitigate performance degradation in low-bit quantization of DETR models. These methods primarily focus on quantizing GEMM operations. For nonlinear activation functions, lookup table (LUT) techniques [46] are commonly used. Additionally, methods like I-BERT [18] and I-ViT [22] employ integer approximation to achieve fixed-point computation.
Quantization for 3D Object Detection. Quantization methods have been applied to accelerate 3D object detection in autonomous driving and robotics. Leveraging advances in image quantization, QD-BEV [66] employs QAT and distillation in multi-camera 3D detection, achieving smaller models and faster inference than the BEVFormer baseline [23]. For LiDAR-based detection, LIDAR-PTQ [68] achieves state-of-the-art quantization on CenterPoint [58], with performance close to FP32 and a threefold speedup. To our knowledge, no PTQ solution currently exists for transformer-based 3D object detection in autonomous driving.
3 Quantization and Deployment-Friendly Adaptation of PETR
In this section, we aim to improve PETR’s quantization performance. We begin by introducing the principles of PETR (§3.1), identify its quantization failures (§3.2), and provide strategies to address these challenges (§3.3).
3.1 Preliminaries
PETR
enhances 2D image features with 3D position-aware properties using camera-ray positional encoding (PE), enabling refined query updates for 3D bounding box prediction. Specifically, surround-view images pass through a backbone to generate 2D features , while camera-ray PE is computed using camera intrinsics and extrinsics. The learnable query embeddings serve as the initial queries for the decoder. Here, serves as the values , and adding to element-wise forms the 3D position-aware keys .
The decoder updates the queries using these key-value pairs through self-attention, cross-attention, and feed-forward network (FFN) modules. The updated query vectors are passed through an MLP to predict 3D bounding box categories and attributes, repeating for cycles. The entire PETR process is summarized in Algorithm 1.
3.2 Quantization Failure of PETR
We evaluate the performance of several PETR configurations [29] using the official code. Under standard 8-bit symmetric per-tensor post-training quantization (PTQ), PETR suffers significant performance degradation, with an average drop of 58.2% in mAP and 36.9% in NDS on the nuScenes validation dataset (see Table 1).
| Bac | Size | Feat | FP32 Acc | INT8 Acc | ||
| mAP | NDS | mAP | NDS | |||
| 0R500 | 1408512 | c5 | 30.5 | 35.0 | 18.4(12.1) | 27.3(07.7) |
| 0R500 | 1408512 | p4 | 31.7 | 36.7 | 15.7(16.0) | 26.1(10.6) |
| V2-99 | 0800320 | p4 | 37.8 | 42.6 | 10.9(26.9) | 23.6(19.0) |
| V2-99 | 1600640 | p4 | 40.4 | 45.5 | 11.3(29.1) | 23.9(21.6) |
Layer-wise Quantization Error Analysis.
Quantizing a pre-trained network introduces output noise, degrading performance. To identify the root causes of quantization failure, we employ the signal-to-quantization-noise ratio (SQNR), inspired by recent PTQ advancements [36, 57, 35]:
Here, is the number of calibration data points; denotes the full-precision network; the quantization error is ; and represents the network output with quantization applied at bit-width , while other layers remain at full precision.
Since 8-bit weight quantization incurs minimal information loss, we focus on quantization errors related to operator inputs. Using the PETR configuration from the first row of Table 1, we obtain layer-wise SQNRs, depicted in Fig. 1. From these results, we identify three primary observations contributing to quantization errors:
Observation 1: The inverse-sigmoid operator disrupts feature distribution balance.
As indicated by the red arrow in Fig.1, quantization difficulties stem from PETR’s positional encoding module. Analyzing its construction (Fig.5(a)), we find that the inverse-sigmoid operation induces an imbalanced feature distribution. Specifically, Fig. 2 shows that before applying inverse-sigmoid, the feature distribution is balanced and quantization-friendly, whereas afterward, it exhibits significant outliers. Thus, avoiding the inverse-sigmoid operator in the positional encoder is advisable.
Observation 2: Disparity in magnitude between camera-ray PE and image features leads to information loss after quantization.
The issue highlighted by the purple arrows in Fig.1 indicates that applying 8-bit symmetric linear quantization to the 3D position-aware key significantly degrades performance. Since is derived by element-wise summing image features and camera-ray PE, we quantize them using ’s quantization parameters, as visualized in Fig.3. Clearly, image features occupy only 4 bits, causing severe information loss. Designing a positional encoding with a magnitude distribution similar to image features is crucial to address this quantization issue.
Observation 3: The scaled dot-product in cross-attention exhibits significant fluctuations along the head dimension.
As highlighted by the green arrow in Fig.1, the scaled dot-product in cross-attention presents quantization challenges. Since the subsequent softmax operation is applied along the last dimension, we visualize the maximum values over this dimension in Fig.4. The feature distribution varies significantly across different heads, posing challenges for effective per-tensor quantization.
3.3 Quantization and Deployment Friendly Improvement.
Based on the analysis in Section 3.2, we identify two key areas for improvement: (1) designing a positional encoding with magnitude and distribution similar to image features, and (2) balancing the scaled dot-products in cross-attention across heads. We propose specific methods to address these challenges.
Positional Encoding Adaptation.
To reduce the magnitude of positional encoding in PETR, we analyze its construction process (Fig.5(a)). PETR samples 64 3D points along each camera ray per pixel, increasing magnitude variance. According to statistical theory, the variance of a sum increases with the number of variables. Inspired by 3DPPE [43], which introduces a lidar-centered lidar-ray PE using only one 3D point (Fig.5(b)), we design a quantization and deployment-friendly (QD-aware) lidar-ray PE (Fig.5(c)).
Our QD-aware lidar-ray PE avoids both the inverse-sigmoid operator and sinusoidal calculations, making it suitable for quantization and deployment on edge devices. Specifically, we learn three anchor embeddings along each axis , denoted with anchor locations for . For each lidar-ray point , we identify the two closest anchors and compute interpolated embeddings:
The QD-aware lidar-ray PE is then obtained via , where denotes concatenation along the channel dimension, and the MLP has a conv-relu-conv structure.
Quantization Strategy for Scaled Dot-Product in Cross-Attention.
In softmax operations, numerical stabilization (NS) subtracts the maximum value to prevent overflow. Traditional quantization quantizes before NS, leading to issues in Observation3. We propose quantizing after NS (Fig.6), and adaptively determining the optimal truncation lower bound to minimize softmax error.
After NS, inputs to softmax are non-positive. Values below approach zero after exponentiation, so we define a candidate set of scaling factors with for -bit quantization. The dequantized input is:
ensuring . We compute the softmax distributions and , and select the optimal scaling factor minimizing the error:
DuLUT for Non-linear Functions.
Lookup table (LUT) is commonly used for integer inference of non-linear functions in quantized neural networks. Linear LUT can compute functions losslessly when the input quantization bit width is less than or equal to the number of LUT entries. However, for higher-bit inputs, conventional linear LUT become impractical due to exponentially increasing table sizes—requiring 256 entries for 8-bit inputs and 65,536 entries for 16-bit inputs—leading to significant storage and computational overhead.
To improve precision without excessive resource usage, non-linear LUT have been proposed. These methods allocate more entries to regions where the function changes rapidly and fewer entries where it remains flat, thus enhancing quantization precision in critical areas. However, non-linear LUT are complex to implement in hardware, necessitating additional processing to define input regions based on function characteristics, and are not efficiently supported by most hardware architectures. Approximation methods like I-BERT [18] and I-ViT [22] attempt to simplify computations but introduce approximation errors and increased computational steps, offsetting the benefits of quantization.
To address these challenges, we propose DuLUT, which uses two small linear tables tailored to the function’s properties, enabling accurate non-linear approximation with significantly reduced table sizes. DuLUT splits the input domain into two regions—typically where the function is rapidly changing and where it is relatively flat—and employs a separate small linear LUT for each region. For example, with 8-bit quantization, DuLUT uses two tables of 32 entries each without compromising precision (see Fig.7 and Algorithm 2). We applied DuLUT to common activation functions like softmax, GELU, and SiLU. By utilizing DuLUT, we achieve the same precision as larger single-table lookups while significantly reducing SRAM overhead and maintaining computational efficiency.
4 Experiment
4.1 Experimental Setup
Benchmark. We use the nuScenes dataset, a comprehensive autonomous driving dataset covering object detection, tracking, and LiDAR segmentation. The vehicle is equipped with one LiDAR, five radars, and six cameras providing a 360-degree view. The dataset comprises 1,000 driving scenes split into training (700 scenes), validation (150 scenes), and testing (150 scenes) subsets. Each scene lasts 20 seconds, annotated at 2 Hz.
Metrics. Following the official evaluation protocol, we report the nuScenes Score (NDS), mean Average Precision (mAP), and five true positive metrics: mean Average Translation Error (mATE), Scale Error (mASE), Orientation Error (mAOE), Velocity Error (mAVE), and Attribute Error (mAAE).
Experimental Details. Our experiments encompass both floating-point training and quantization configurations. For floating-point training, we follow PETR series settings, using PETR with an R50dcn backbone unless specified, and utilize the C5 feature (1/32 resolution output) as the 2D feature. Input images are at resolution. Both the lidar-ray PE and QD-aware lidar-ray PE use a pixel-wise depth of 30m with three anchor embeddings per axis. The 3D perception space is defined as m along the X and Y axes, and m along the Z axis. We also compare these positional encodings on StreamPETR, using a V2-99 backbone and input images of resolution.
Training uses the AdamW optimizer (weight decay 0.01) with an initial learning rate of , decayed via a cosine annealing schedule. We train for 24 epochs with a batch size of 8 on four NVIDIA RTX 4090 GPUs. No test-time augmentation is applied.
For quantization, we adopt 8-bit symmetric per-tensor post-training quantization, using 32 randomly selected training images for calibration. When quantizing the scaled dot-product in cross-attention, we define a candidate set of 20 scaling factors.
4.2 Ablation Study
Proof of Position Encoding Equivalence We conducted experiments to verify whether the proposed QD-aware lidar-ray PE enhances floating-point performance over the original camera-ray PE. As shown in Tab. 2, QD-aware lidar-ray PE provides performance improvements. On PETR, it slightly increases mAP by 0.07 but significantly boosts NDS and mATE by 1.09 and 1.67, respectively. For Stream-PETR, our method yields substantial and balanced enhancements, with increases of 0.94 in mAP, 0.46 in NDS, and 0.22 in mATE.
| Method | mAP | NDS | mATE | |
| PETR | Camera-ray PE | 31.42(0.00) | 36.11(0.00) | 84.19(0.00) |
| QD-aware lidar-ray PE | 31.49(0.07) | 37.20(1.09) | 82.52(1.67) | |
| Stream -PETR | Camera-ray PE | 48.19(0.00) | 57.11(0.00) | 60.99(0.00) |
| QD-aware lidar-ray PE | 49.13(0.94) | 57.57(0.46) | 60.77(0.22) | |
Effect of Anchor Embedding Quantity. The QD-aware lidar-ray PE uses three anchor embeddings per axis, obtained through linear interpolation. Experiments (Tab. 3) demonstrate that setting the number of anchor embeddings to 3 achieves the highest NDS and mAP scores. Adjusting this number either up or down results in lower performance, confirming that 3 is the optimal choice.
| Quantity of Anchor Embedding | NDS | mAP | ||
| x-axis | y-axis | z-axis | ||
| 2 | 2 | 2 | 36.66 | 31.29 |
| 3 | 3 | 3 | 37.20 | 31.49 |
| 4 | 4 | 4 | 36.92 | 31.09 |
| 5 | 5 | 5 | 36.83 | 31.19 |
Quantization Performance of Different Position Encodings To experimentally demonstrate the superior quantization performance of our proposed QD-aware lidar-ray PE, we focus solely on quantizing the positional encoding, keeping all other modules in floating-point computation. Detailed results are shown in Tab.4. The original camera-ray configuration loses up to 11.97% in mAP and 5.04% in NDS, whereas our QD-aware lidar-ray PE experiences minimal losses of only 1.42% in mAP and 1.15% in NDS. Figure8 further supports this finding; compared to the distribution in Fig. 3, the distribution of our QD-aware lidar-ray PE aligns more closely with that of image features, retaining sufficient useful information.
| Method | NDS | mAP | mATE | |
| PETR | Camera-ray PE | 34.29 | 27.66 | 87.17 |
| QD-aware lidar-ray PE | 36.77 | 31.04 | 82.92 | |
| Stream -PETR | Camera-ray PE | 53.74 | 40.23 | 69.39 |
| QD-aware lidar-ray PE | 56.81 | 47.65 | 61.53 | |
Impact of Different Scaled Dot Product Quantization Strategy. To validate our novel scaled dot-product quantization strategy—which searches for the optimal scaling factor by minimizing softmax output error—we focus solely on quantizing the softmax inputs while keeping all other modules in floating-point computation. As shown in Tab. 5, the original quantization strategy results in significant losses of 40% in NDS and 50% in mAP. In contrast, our ”quant after stabilization” approach greatly improves performance. An ablation study on the maximum candidate truncation range reveals that setting yields optimal quantization performance with nearly no loss. Performance deteriorates when due to increased truncation of feature information, while values of exceeding 20 offer no additional benefits. Therefore, setting is sufficient to achieve the best performance.
| Method | NDS | mAP | mATE | |
| quant before stabilization | - | 25.31 | 13.79 | 107.94 |
| quant after stabilization | N = 10 | 3.45 | 1.23 | 150.34 |
| N = 50 | 33.86 | 28.77 | 87.12 | |
| N = 10 | 34.65 | 29.33 | 85.01 | |
| N = 20 | 36.10 | 31.42 | 84.19 | |
| N = 30 | 36.10 | 31.42 | 84.19 | |
| N = 40 | 36.10 | 31.42 | 84.19 | |
| Backbone | Input Size | Feature | Mode | mAP | NDS | mATE | mASE | mAOE | mAVE | mAAE | |
| R50 | 5121408 | c5 | PETR | FP32 | 31.42(0.00) | 36.11(0.00) | 84.19(0.00) | 28.42(0.00) | 60.69(0.00) | 99.08(0.00) | 23.58(0.00) |
| PTQ | 13.79[56.11] | 25.31[29.91] | 107.94[28.21] | 31.47[10.73] | 75.22[23.94] | 83.71[15.51] | 25.45[7.93] | ||||
| Q-PETR | FP32 | 31.49(0.07) | 37.20(1.09) | 82.52(1.67) | 27.88(0.54) | 59.91(0.78) | 91.74(7.34) | 23.45(0.13) | |||
| PTQ | 31.34[0.47] | 37.17[0.82] | 82.61[0.65] | 27.93[0.17] | 60.00[0.15] | 91.79[0.00] | 23.45[0.00] | ||||
| R50 | 5121408 | p4 | PETR | FP32 | 32.60(0.00) | 37.16(0.00) | 82.63(0.00) | 27.96(0.00) | 61.06(0.00) | 95.81(0.00) | 23.91(0.00) |
| PTQ | 12.97[60.21] | 24.75[33.39] | 108.28[31.04] | 31.76[13.59] | 79.57[30.31] | 78.90[17.65] | 27.14[13.51] | ||||
| Q-PETR | FP32 | 32.69(0.09) | 38.03(0.87) | 80.58(2.05) | 27.89(0.07) | 59.43(0.63) | 92.69(3.12) | 22.55(1.36) | |||
| PTQ | 32.40[0.88] | 37.72[0.82] | 81.11[0.65] | 27.92[0.10] | 60.02[0.97] | 92.76[0.00] | 22.59[0.18] | ||||
| R101 | 5121408 | p4 | PETR | FP32 | 34.40(0.00) | 38.62(0.00) | 80.67(0.00) | 28.03(0.00) | 57.13(0.00) | 95.74(0.00) | 24.20(0.00) |
| PTQ | 13.53[60.67] | 23.84[38.27] | 111.04[37.65] | 31.27[11.56] | 78.94[38.18] | 92.92[2.95] | 26.14[8.02] | ||||
| Q-PETR | FP32 | 34.72(0.32) | 39.68(1.06) | 79.40(1.27) | 27.92(0.11) | 53.90(3.23) | 92.99(2.75) | 22.59(1.61) | |||
| PTQ | 34.41[0.89] | 39.08[1.51] | 80.98[1.98] | 28.00[0.28] | 54.41[0.94] | 92.62[0.39] | 22.70[0.47] | ||||
| V2-99 | 320800 | p4 | PETR | FP32 | 38.01(0.00) | 42.56(0.00) | 75.79(0.00) | 26.84(0.00) | 50.58(0.00) | 90.13(0.00) | 21.07(0.00) |
| PTQ | 10.46[72.48] | 23.64[44.45] | 112.41[36.21] | 33.00[22.95] | 85.96[69.95] | 71.83[20.30] | 25.12[19.22] | ||||
| Q-PETR | FP32 | 38.43(0.42) | 43.80(1.24) | 74.79(1.00) | 27.29(0.45) | 49.76(0.82) | 82.11(8.02) | 20.15(0.92) | |||
| PTQ | 37.93[1.30] | 43.17[1.44] | 75.50[0.94] | 27.78[1.79] | 50.07[0.62] | 82.41[0.36] | 20.38[1.14] | ||||
| V2-99 | 6401600 | p4 | PETR | FP32 | 40.66(0.00) | 46.05(0.00) | 71.76(0.00) | 27.07(0.00) | 42.23(0.00) | 80.68(0.00) | 21.06(0.00) |
| PTQ | 6.40[84.63] | 20.98[54.55] | 117.38[33.22] | 34.85[25.92] | 83.38[97.44] | 76.83[4.79] | 27.10[28.67] | ||||
| Q-PETR | FP32 | 41.35(0.69) | 45.99(0.06) | 72.18(0.42) | 26.91(0.15) | 45.05(2.82) | 82.03(2.35) | 20.67(0.39) | |||
| PTQ | 40.95[0.96] | 45.64[1.09] | 73.40[1.69] | 27.05[0.52] | 45.61[1.24] | 82.17[0.17] | 20.68[0.00] | ||||
| V2-99 | 320800 | p4 | StreamPETR | FP32 | 48.19(0.00) | 57.11(0.00) | 60.99(0.00) | 25.58(0.00) | 37.54(0.00) | 26.28(0.00) | 19.43(0.00) |
| PTQ | 18.52(61.19) | 36.47(36.11) | 76.39(25.25) | 31.44 (22.90) | 47.03 (25.27) | 30.22 14.99 | 20.99 (8.03) | ||||
| Q-StreamPETR | FP32 | 49.13(0.94) | 57.57(0.46) | 60.77(0.22) | 26.14(0.56) | 39.05(1.49) | 24.81(1.47) | 19.15(0.28) | |||
| PTQ | 48.21 (1.87) | 56.33 (2.15) | 63.00 (3.66) | 26.35 (0.80) | 39.17 (0.31) | 24.90 (0.36) | 19.19 (0.21) | ||||
| V2-99 | 6401600 | p4 | StreamPETR | FP32 | 49.51(0.00) | 58.03(0.00) | 60.10(0.00) | 26.07(0.00) | 35.65(0.00) | 25.91(0.00) | 19.60(0.00) |
| PTQ | 18.72(62.19) | 35.66(38.54) | 74.32(23.66) | 30.39 (16.76) | 41.49 (16.38) | 30.53 17.83 | 20.82 (6.22) | ||||
| Q-StreamPETR | FP32 | 50.48(0.93) | 58.61(0.58) | 58.78(0.32) | 26.16(0.09) | 37.05(1.40) | 25.69(0.22) | 18.59(1.01) | |||
| PTQ | 49.44 (2.06) | 57.94 (1.24) | 60.30 (2.52) | 26.55 (1.49) | 37.07 (0.00) | 25.87 (0.31) | 18.59 (0.00) | ||||
Superior Performance of DuLUT for Non-linear Functions. To validate the quantization advantages of our proposed DuLUT for nonlinear functions, we use ”quant after stabilization (N=20)” from Tab. 5 as a baseline and evaluate the performance with different nonlinear function quantization methods applied on top of it. The specific results are shown in Tab. 7. We consider the carefully designed approximation methods I-Bert and I-Vit for different nonlinear functions. Due to the approximation errors introduced by these methods, many points are quantized away. Additionally, we compare with the LUT-based table lookup method and find that 256 entries are required for lossless quantization, while 128 entries lead to severe performance losses of 0.54 NDS and 0.37 mAP. In contrast, our newly proposed DuLUT with 128 entries achieves lossless quantization. Even when the number of entries is further reduced to 64, the quantization only results in a negligible loss of 0.08% NDS and 0.02% mAP, which can be considered negligible. This further demonstrates the superior quantization performance of our proposed DuLUT.
| Method | NDS | mAP | mATE | |
| Quant after stabilization (N=20) | 36.10 | 31.42 | 84.19 | |
| I-Bert | 34.87 | 29.34 | 88.41 | |
| I-Vit | 35.03 | 28.77 | 87.32 | |
| LUT | 256 entries | 36.10 | 31.42 | 84.19 |
| 128 entries | 35.56 | 31.05 | 85.61 | |
| DuLUT | (16,16) entries | 28.12 | 17.36 | 96.99 |
| (16,32) entries | 34.14 | 27.29 | 90.33 | |
| (32,32) entries | 36.07 | 31.36 | 84.20 | |
| (64,64) entries | 36.10 | 31.42 | 84.19 | |
4.3 Validation on PETR-series Methods
We evaluate the effectiveness of our proposed method on various PETR-series models from both floating-point and quantized performance perspectives, specifically considering single-frame PETR and temporal multi-frame StreamPETR models.
First, we analyze changes in floating-point performance (values in parentheses in Tab. 6). In single-frame PETR models, mAP and NDS generally improve across configurations, except for a slight decrease of 0.06 in NDS when using V2-99’s P4 feature with 640×1600 resolution images. mAP increases range from 0.07 to 0.69, while NDS shows significant gains in all cases, ranging from 0.87 to 1.24. For temporal multi-frame StreamPETR models, both mAP and NDS consistently improve, with mAP gains of 0.93 and 0.94, and modest NDS increases of 0.46 and 0.58. Notably, NDS improvements in temporal methods are smaller than in single-frame methods, mainly due to performance degradation in mASE and mAOE, suggesting that our QD-aware lidar-ray PE may not optimally capture scale and orientation information in temporal models. We plan to investigate this further in future work. Overall, QPETR shows significant improvements in most configuration metrics, demonstrating that our method surpasses the original PETR models in floating-point performance.
Next, we analyze the quantized performance improvements (values in square brackets in Table 6). In single-frame PETR models, mAP and NDS drops are kept below 1% using our QD-aware lidar-ray PE and smoothing techniques. In temporal multi-frame StreamPETR models, mAP and NDS drops remain within 2.5%, likely due to accumulated quantization errors during temporal fusion. Overall, quantized QPETR models maintain high performance with minimal drops in both settings, demonstrating the effectiveness of our quantization strategies in preserving accuracy while reducing computational and memory requirements. We intend to further mitigate quantization errors in temporal models through enhanced error correction or advanced quantization methods.
5 Limitations
Although our method incurs almost no quantization accuracy loss, users need to replace the camera-ray in the original PETR series with our proposed QD-aware lidar-ray. The only drawback is that this requires retraining. However, from the perspective of quantization deployment, this retraining is beneficial, and the floating-point precision can even be improved.
6 Conclusion
In this paper, we address the significant performance drops of PETR models during quantization by identifying two main issues: the imbalance between positional encoding and image feature magnitudes, and uneven scalar dot-products in cross-attention. To resolve these, we introduce Q-PETR, a quantization-friendly positional encoding transformation that redesigns positional encoding and improves scalar dot-product quantization without sacrificing the original floating-point performance. Our experiments show that Q-PETR limits mAP and NDS drops to below 1% under standard 8-bit post-training quantization and even surpasses the original PETR in floating-point precision. Extensive tests across various PETR models demonstrate the method’s strong generalization and deployment suitability. Future work will explore quantization error propagation in multi-frame temporal models and develop advanced quantization strategies to further reduce precision loss in temporal settings.

### v2

@@
-Q-PETR: Quant-aware Position Embedding Transformation for Multi-View 3D Object Detection
+Q-PETR: Quantization-aware Position Embedding Transformation for Multi-View 3D Object Detection
 Abstract
-PETR-based methods have dominated benchmarks in 3D perception and are increasingly becoming a key component in modern autonomous driving systems.
-However, their quantization performance significantly degrades when INT8 inference is required, with a degradation of 58.2% in mAP and 36.9% in NDS on the NuScenes dataset.
-To address this issue, we propose a quantization-aware position embedding transformation for multi-view 3D object detection, termed Q-PETR.
-Q-PETR offers a quantization-friendly and deployment-friendly architecture while preserving the original performance of PETR.
-It substantially narrows the accuracy gap between INT8 and FP32 inference for PETR-series methods.
-Without bells and whistles, our approach reduces the mAP and NDS drop to within 1% under standard 8-bit per-tensor post-training quantization.
-Furthermore, our method exceeds the performance of the original PETR in terms of floating-point precision.
-Extensive experiments across a variety of PETR-series models demonstrate its broad generalization.
+Camera-based multi-view 3D detection has emerged as an attractive solution for autonomous driving due to its low cost and broad applicability.
+However, despite the strong performance of PETR-based methods in 3D perception benchmarks, their direct INT8 quantization for onboard deployment leads to drastic accuracy drops—up to 58.2% in mAP and 36.9% in NDS on the NuScenes dataset.
+In this work, we propose Q-PETR, a quantization-aware position embedding transformation that re-engineers key components of the PETR framework to reconcile the discrepancy between the dynamic ranges of positional encodings and image features, and to adapt the cross-attention mechanism for low-bit inference.
+By redesigning the positional encoding module and introducing an adaptive quantization strategy, Q-PETR maintains floating-point performance with a performance degradation of less than 1% under standard 8-bit per-tensor post-training quantization.
+Moreover, compared to its FP32 counterpart, Q-PETR achieves a two-fold speedup and reduces memory usage by three times, thereby offering a deployment-friendly solution for resource-constrained onboard devices.
+Extensive experiments across various PETR-series models validate the strong generalization and practical benefits of our approach.
 1 Introduction
-3D perception is crucial for autonomous driving and embodied intelligence [17, 67, 31, 16, 62].
-PETR-based methods [49, 29, 30, 43, 56, 50, 52] have gained significant attention due to their effectiveness in this domain.
-Unlike previous detection methods relying on dense convolutional features [17, 61, 60, 59, 37] or utilizing 3D-to-2D projections as in DETR3D-based approaches [51, 25, 28], PETR adapts the paradigm of 2D object detection with transformers (DETR) [5] by incorporating 3D positional encoding, achieving seamless end-to-end 3D obstacle detection.
-Despite its advantages, PETR requires substantial storage and computational resources during inference, posing challenges for on-device deployment.
-To enhance inference efficiency, considerable efforts have been dedicated to efficient model compression [42, 20, 27, 19, 11, 12], with quantization being particularly favored for deploying models on AI chips by representing networks in low-bit formats [39, 32, 54, 7].
-However, deploying complex models like PETR on edge devices presents additional challenges.
-Autonomous vehicles rely on edge AI chips with limited computational capabilities.
-Nonlinear functions such as softmax, sigmoid, SiLU, and GeLU are commonly used during inference, but most edge AI chips lack sufficient floating-point performance for these operations, exacerbating deployment difficulties.
-Moreover, quantization awareness is often overlooked in neural architecture design, leading to significant performance degradation during deployment.
-Notably, models like MobileNet [40, 64, 15, 38], EfficientNet [4, 13, 45], and RepVGG [8] face quantization issues requiring additional design efforts or advanced strategies to mitigate deployment challenges.
-PETR is no exception; when deployed on in-vehicle AI chips, standard 8-bit per-tensor post-training quantization (PTQ) results in a significant performance drop, with a 58.2% decrease in mAP and a 36.9% decrease in NDS.
-In this paper, we address the quantization issues of PETR to facilitate its deployment on edge AI chips in autonomous vehicles.
-We conduct an in-depth analysis of the mechanisms behind PETR’s quantization failure, revealing that the significantly larger magnitudes of positional encodings compared to image features, and the highly imbalanced scaled dot-products in cross-attention, are the primary causes of quantization collapse.
-Based on these insights, we propose Q-PETR (Quantization-aware PETR), a model that mitigates quantization collapse and even improves floating-point performance compared to the original PETR.
-Our contributions are summarized as follows:
-Identifying the root causes of quantization collapse in PETR: We uncover that the imbalanced output of the inverse-sigmoid function and excessively large values in positional encoding, as well as the imbalanced scaled dot-products in cross-attention, are key issues.
-Redesigning the positional encoding and quantization strategy: We reformulate the positional encoding and improve the quantization approach for the scaled dot-product in cross-attention, enhancing both floating-point performance and quantization friendliness.
-Introducing DuLUT for efficient nonlinear function inference: We propose DuLUT, an improved lookup table [46] approach that reduces the number of table entries without compromising precision, optimizing nonlinear function inference on edge AI chips.
-Demonstrating generalizability and deployment readiness: Our method generalizes well across different model scales, achieving exceptional post-quantization performance suitable for deployment on edge devices.
+3D object detection [58, 59, 61, 26] has been a longstanding topic in computer vision.
+Compared to LiDAR, cameras have gained increasing popularity in autonomous systems due to their ability to provide dense texture information at a lower cost, making camera-based 3D object detection more favored [13, 12, 54].
+Among these studies, the mainstream PETR frameworks [39, 23, 24, 34, 47, 40, 42] have gained prominence by adapting the 2D transformer-based DETR paradigm [5] with 3D positional encodings.
+In comparison to dense feature methods [13, 53, 52, 50, 32] or DETR3D-style 3D-to-2D projections [41, 19, 21], PETR achieves end-to-end 3D detection while performing superior performance.
+Despite its effective, the deployment of PETR on resource-limited edge devices presents a critical challenge that significantly hinders their widespread application in autonomous vehicles and robotics.
+Quantization [60, 57, 7, 29] is an efficient model compression approach that reduces computational burden by converting high-bit floating-point into low-bit integer formats.
+Compared to quantization-aware training (QAT) methods, which require access to all labeled training data and substantial computation resources, post-training quantization (PTQ) is more suitable for rapid and efficient paratical applications.
+This is because PTQ only needs a small number of unlabeled samples for calibration.
+Furthermore, PTQ does not require retraining the network with all available labeled dataset, resulting in a shorter implementation time.
+Although several advanced PTQ methods have been proposed for 2D detection tasks[15, 43, 22] and ViT [16, 25], directly applying them to multi-view 3D Detection tasks inevitably leads to severe performance degradation due to the structures and task-specific differences.
+For instance, when standard 8-bit per-tensor post-training quantization (PTQ) is applied to PETR, leading up to 58.2% mAP and 36.9% NDS performance collapse.
+Furthremore, nonlinear operators such as softmax, gelu, and silu are indispensable in 3D detection models, but their performance is often hindered by hardware constraints, akin to the “short board” in a barrel.
+Even specialized Tensor Cores in high-end GPUs like the NVIDIA A100 exhibit significantly lower throughput for these nonlinear operators compared to matrix multiplications [9].
+Moreover, many edge AI chips rely on a lookup table (LUT) [36] for integer activation functions, but these typically only support integer inputs and are often linear LUTs.
+While a linear LUT can faithfully approximate nonlinearities if its size covers the entire dynamic range, its capacity grows exponentially (e.g., from 256 entries at 8-bit to 65536 entries at 16-bit), making it impractical.
+Non-linear LUTs [51] allocate more entries to steeper regions of the function and fewer to flatter regions, but introduce additional hardware complexity.
+Other integer-only methods [14, 16] further simplify operator emulation at the cost of increased approximation error.
+Consequently, effective quantization and hardware-friendly acceleration of these nonlinear operators is crucial and unexplored for PETR in resource-constrained deployment.
+In this paper, we address these quantization challenges for PETR-based 3D Detection.
+Firstly, through an in-depth analysis, we find that disproportionately large positional encodings and imbalanced scaled dot-products in cross-attention significantly degrade quantized performance.
+Building on these findings, we propose Q-PETR, a quantization-friendly variant of PETR that not only mitigates performance collapse but also enhances floating-point accuracy.
+To further resolve the bottleneck of nonlinear operators, we introduce a lightweight dual-LUT (DuLUT) mechanism that maintains high approximation fidelity with fewer table entries.
+Our main contributions are summarized as follows:
+Diagnosis of quantization failures in PETR: We show that large positional encodings, imbalanced inverse-sigmoid outputs, and skewed cross-attention dot-products are key factors causing significant accuracy loss under low-bit quantization.
+Redesign of positional encodings and cross-attention quantization: We reformulate the positional encoding module and employ a more balanced scaling strategy for cross-attention dot-products, improving both floating-point and quantized performance.
+Introduction of DuLUT for hardware-friendly nonlinear functions: By splitting LUT entries based on the curvature of the function, DuLUT efficiently approximates nonlinearities with fewer table entries, greatly facilitating edge deployment.
+Broad applicability and deployment readiness: Our approach generalizes to various model scales, achieving minimal performance loss under standard 8-bit PTQ while even boosting full-precision accuracy, thus meeting the demands of resource-limited scenarios in real-world.
 2 Related Work
-3D Object Detection.
+Multi-View 3D Object Detection.
 Surround-view 3D object detection is essential for autonomous driving and is generally categorized into LSS-based [13, 50, 12] and transformer-based [23, 34] approaches.
@@
 QAT [10, 4] introduces quantization-aware losses during training, enhancing robustness but requiring resource-intensive retraining.
-PTQ offers rapid deployment without retraining.
-While PTQ methods have been successful on CNNs [33, 34, 21], they often perform poorly on Vision Transformers (ViTs) and transformer-based detectors due to structural differences.
-For ViTs, practical PTQ algorithms have been developed [63, 26, 44, 24, 41].
+Compare to QAT, PTQ offers rapid deployment without retraining.
+While PTQ methods have been successful on CNNs [28, 29, 15], they often perform poorly on transformer-based 3D detectors due to structural differences.
+For ViTs, practical PTQ algorithms have been developed [55, 20, 35, 18].
 In the context of transformer-based object detection models, Q-DETR [46] and AQ-DETR [37] use QAT and knowledge distillation to mitigate performance degradation in low-bit quantization of DETR models.
@@
 Leveraging advances in image quantization, QD-BEV [57] employs QAT and distillation in multi-camera 3D detection, achieving smaller models and faster inference than the BEVFormer baseline [17].
-For LiDAR-based detection, LIDAR-PTQ [68] achieves state-of-the-art quantization on CenterPoint [58], with performance close to FP32 and a threefold speedup.
-To our knowledge, no PTQ solution currently exists for transformer-based 3D object detection in autonomous driving.
+For LiDAR-based detection, LIDAR-PTQ [60] achieves state-of-the-art quantization on CenterPoint [49], with performance close to FP32 and 3 speedup.
+To our knowledge, there are no PTQ solution tailored for transformer-based 3D detection in autonomous driving.
 3 Quantization and Deployment-Friendly Adaptation of PETR
 In this section, we aim to improve PETR’s quantization performance.
-We begin by introducing the principles of PETR (§3.1), identify its quantization failures (§3.2), and provide strategies to address these challenges (§3.3).
-3.1 Preliminaries
-PETR
-enhances 2D image features with 3D position-aware properties using camera-ray positional encoding (PE), enabling refined query updates for 3D bounding box prediction.
-Specifically, surround-view images pass through a backbone to generate 2D features , while camera-ray PE is computed using camera intrinsics and extrinsics.
-The learnable query embeddings serve as the initial queries for the decoder.
-Here, serves as the values , and adding to element-wise forms the 3D position-aware keys .
-The decoder updates the queries using these key-value pairs through self-attention, cross-attention, and feed-forward network (FFN) modules.
-The updated query vectors are passed through an MLP to predict 3D bounding box categories and attributes, repeating for cycles.
-The entire PETR process is summarized in Algorithm 1.
+We begin by elaborating the principles of PETR in Sup.
+A, identify its quantization failures (§3.1), and provide strategies to address these challenges (§3.2).
 3.1 Quantization Failure of PETR
@@
 To identify the root causes of quantization failure, we employ the signal-to-quantization-noise ratio (SQNR), inspired by recent PTQ advancements [31, 48, 30]:
-Here, is the number of calibration data points; denotes the full-precision network; the quantization error is ; and represents the network output with quantization applied at bit-width , while other layers remain at full precision.
-Since 8-bit weight quantization incurs minimal information loss, we focus on quantization errors related to operator inputs.
+Here, is the number of calibration data points; denotes the full-precision network; the quantization error is ; and denotes the network output when only the target layer is quantized to bits, with all other layers kept at full precision.
+Since 8-bit weight quantization results in only a minor loss of precision, we focus on quantization errors arising from operator inputs.
 Using the PETR configuration from the first row of Table 1, we obtain layer-wise SQNRs, depicted in Fig.
 1.
-From these results, we identify three primary observations contributing to quantization errors:
-Observation 1: The inverse-sigmoid operator disrupts feature distribution balance.
-As indicated by the red arrow in Fig.1, quantization difficulties stem from PETR’s positional encoding module.
-Analyzing its construction (Fig.5(a)), we find that the inverse-sigmoid operation induces an imbalanced feature distribution.
-Specifically, Fig.
-2 shows that before applying inverse-sigmoid, the feature distribution is balanced and quantization-friendly, whereas afterward, it exhibits significant outliers.
-Thus, avoiding the inverse-sigmoid operator in the positional encoder is advisable.
-Observation 2: Disparity in magnitude between camera-ray PE and image features leads to information loss after quantization.
-The issue highlighted by the purple arrows in Fig.1 indicates that applying 8-bit symmetric linear quantization to the 3D position-aware key significantly degrades performance.
-Since is derived by element-wise summing image features and camera-ray PE, we quantize them using ’s quantization parameters, as visualized in Fig.3.
-Clearly, image features occupy only 4 bits, causing severe information loss.
-Designing a positional encoding with a magnitude distribution similar to image features is crucial to address this quantization issue.
-Observation 3: The scaled dot-product in cross-attention exhibits significant fluctuations along the head dimension.
-As highlighted by the green arrow in Fig.1, the scaled dot-product in cross-attention presents quantization challenges.
-Since the subsequent softmax operation is applied along the last dimension, we visualize the maximum values over this dimension in Fig.4.
-The feature distribution varies significantly across different heads, posing challenges for effective per-tensor quantization.
+From these results, we identify three main factors contributing to quantization errors:
+Observation 1: Position Encoding Design Flaws Lead to Quantization Difficulties.
+We find that PETR’s quantization issues primarily arise from its positional encoding module in two key ways.
+(a) Inverse-sigmoid disrupts feature balance.
+As shown in Fig.
+1, the inverse-sigmoid operation skews an otherwise balanced distribution (Fig.
+2) toward significant outliers.
+(b) Magnitude disparity between camera-ray PE and image features.
+As highlighted by the purple arrow in Fig.
+1, applying 8-bit quantization to the 3D position-aware key yields severe performance drops.
+Statistical analysis (Fig.
+3) shows that camera-ray PE spans approximately , while image features remain within .
+Consequently, when using Eq.
+2 and an 8-bit symmetric range of , PE dominates the scaling factor.
+Image features then collapse into merely seven bins (Fig.
+4), causing catastrophic information loss and sharp accuracy degradation (Table 1).
+To address these flaws, we (1) remove the inverse-sigmoid step that drives outlier magnitudes, and (2) redesign the positional encoding to align its scale with that of image features.
+This balanced approach preserves critical information during quantization.
+Observation 2: Dual-Dimensional Heterogeneity in Cross-Attention Leads to Quantization Bottlenecks.
+As evidenced by the green arrow in Fig.
+1 and further clarified in Fig.
+5, the scaled dot-product in cross-attention exhibits pronounced heterogeneity on two levels.
+First, the inter-head variance spans 2–3 orders of magnitude, while within each head, the value distribution is extremely broad (e.g., ranging beyond [, ]).
+We merge the head and query dimensions to directly reveal the row-wise feature distribution.
+The results show that regardless of whether quantization is performed per head, per token, or on the entire tensor, the excessively large softmax inputs result in significant quantization errors.
+This confirms that existing quantization paradigms are fundamentally inadequate for handling the severe amplitude disparities in the cross-attention mechanism.
 3.2 Quantization and Deployment Friendly Improvement.
-Based on the analysis in Section 3.2, we identify two key areas for improvement: (1) designing a positional encoding with magnitude and distribution similar to image features, and (2) balancing the scaled dot-products in cross-attention across heads.
-We propose specific methods to address these challenges.
+Drawing on the analysis in Section 3.1 and the deployment challenges noted in the Introduction, we identify three critical issues.
+Firstly, the positional encoding module mismatches the magnitude and distribution of image features, causing severe quantization loss.
+Secondly, imbalanced scaled dot-products in cross-attention further compound quantization errors.
+Thirdly, the high computational cost of nonlinear functions impedes efficient edge inference.
+We propose targeted solutions for above challenges.
 Positional Encoding Adaptation.
-To reduce the magnitude of positional encoding in PETR, we analyze its construction process (Fig.5(a)).
-PETR samples 64 3D points along each camera ray per pixel, increasing magnitude variance.
-According to statistical theory, the variance of a sum increases with the number of variables.
-Inspired by 3DPPE [43], which introduces a lidar-centered lidar-ray PE using only one 3D point (Fig.5(b)), we design a quantization and deployment-friendly (QD-aware) lidar-ray PE (Fig.5(c)).
-Our QD-aware lidar-ray PE avoids both the inverse-sigmoid operator and sinusoidal calculations, making it suitable for quantization and deployment on edge devices.
-Specifically, we learn three anchor embeddings along each axis , denoted with anchor locations for .
-For each lidar-ray point , we identify the two closest anchors and compute interpolated embeddings:
-The QD-aware lidar-ray PE is then obtained via , where denotes concatenation along the channel dimension, and the MLP has a conv-relu-conv structure.
+From the derivations in Appendix C.2, we establish that the amplitude of camera-ray PE can theoretically reach up to 11.5 times that of its LiDAR-based counterpart.
+This stark discrepancy directly explains why PETR’s camera-ray encoding often overshoots the dynamic range of image features, thereby hampering quantization.
+Although LiDAR-ray PE alleviates the amplitude issue, its reliance on high-frequency sinusoidal functions remains problematic for low-bit deployments on edge devices.
+To overcome both amplitude and implementation obstacles, we propose a quantization-deployable LiDAR-ray position embedding (QDPE) that sharply curtails magnitudes while avoiding complex nonlinearities.
+Our design contains two main modifications:
+1.
+(Single-point sampling via LiDAR prior).
+Drawing inspiration from the physical properties of LiDAR sensors, we sample only one 3D point per pixel along each depth ray (Fig.
+6 (b)), in contrast to the multi-sample scheme in PETR’s camera-ray PE.
+By discarding the iterative inverse-sigmoid and sinusoidal transformations, we reduce the overall encoding variance.
+2.
+(Anchor-based bounded embedding with convex-combination constraints).
+As illustrated in Fig.
+6 (c), we learn three axis-aligned anchor embeddings for each spatial axis , with corresponding anchor locations .
+For a LiDAR-sampled 3D point , we compute the embedding along each axis by linear interpolation between the nearest two anchors:
+(4) Theorem 1 guarantees that each axis-wise component is strictly confined within the convex hull of its adjacent anchors.
+We concatenate the three axis-wise embeddings and feed them to a lightweight MLP to obtain the final positional encoding vector.
+These two innovations ensure our QD-aware LiDAR-ray PE remains both bounded in magnitude and free of difficult-to-quantize nonlinearities.
+Fig.
+3 and Fig.
+9 visually demonstrate the elimination of outliers.
+Further, the dynamic range of our QD-aware encoding () is only marginally wider than that of standard image features ()—in stark contrast to PETR’s original ().
+The proposed design eliminates complex nonlinear operations (inverse-sigmoid) and spectral components (high-frequency sinusoids), achieving hardware-compatible computation without compromising geometric fidelity.
 Quantization Strategy for Scaled Dot-Product in Cross-Attention.
 In softmax operations, numerical stabilization (NS) subtracts the maximum value to prevent overflow.
-Traditional quantization quantizes before NS, leading to issues in Observation3.
-We propose quantizing after NS (Fig.6), and adaptively determining the optimal truncation lower bound to minimize softmax error.
-After NS, inputs to softmax are non-positive. Values below approach zero after exponentiation, so we define a candidate set of scaling factors with for -bit quantization. The dequantized input is:
+Traditional quantization quantizes before NS, leading to issues in Observation2.
+We propose quantizing after NS (Fig.
+7), and adaptively determining the optimal truncation lower bound to minimize softmax error.
+After NS, inputs for softmax are non-positive. Values below approach zero after exponentiation, so we define a candidate set of scaling factors with for -bit quantization. The dequantized input is:
 ensuring . We compute the softmax distributions and , and select the optimal scaling factor minimizing the error:
 DuLUT for Non-linear Functions.
-Lookup table (LUT) is commonly used for integer inference of non-linear functions in quantized neural networks.
-Linear LUT can compute functions losslessly when the input quantization bit width is less than or equal to the number of LUT entries.
-However, for higher-bit inputs, conventional linear LUT become impractical due to exponentially increasing table sizes—requiring 256 entries for 8-bit inputs and 65,536 entries for 16-bit inputs—leading to significant storage and computational overhead.
-To improve precision without excessive resource usage, non-linear LUT have been proposed.
-These methods allocate more entries to regions where the function changes rapidly and fewer entries where it remains flat, thus enhancing quantization precision in critical areas.
-However, non-linear LUT are complex to implement in hardware, necessitating additional processing to define input regions based on function characteristics, and are not efficiently supported by most hardware architectures.
-Approximation methods like I-BERT [18] and I-ViT [22] attempt to simplify computations but introduce approximation errors and increased computational steps, offsetting the benefits of quantization.
-To address these challenges, we propose DuLUT, which uses two small linear tables tailored to the function’s properties, enabling accurate non-linear approximation with significantly reduced table sizes.
-DuLUT splits the input domain into two regions—typically where the function is rapidly changing and where it is relatively flat—and employs a separate small linear LUT for each region.
-For example, with 8-bit quantization, DuLUT uses two tables of 32 entries each without compromising precision (see Fig.7 and Algorithm 2).
+The error bound of linear LUT can be formally expressed by the maximum interpolation error theorem.
+Given a twice-differentiable function over interval , the maximum approximation error using linear interpolation satisfies:
+where represents the curvature.
+This result indicates that if the second derivative (i.e., curvature) is large within a sub-interval, must be shortened to control the approximation error.
+Conversely, if the curvature is small, the sub-interval can be lengthened.
+Consequently, more interpolation points (LUT entries) should be assigned where the function changes rapidly, while flatter or near-saturated regions may be merged into fewer intervals.
+Building on this principle, DuLUT partitions the input domain into three types of sub-intervals—shrink (near-saturated), enlarge (steep change), and unchange (near-linear)—defined as follows:
+shrink: for regions where the function is close to saturation or changes very little, multiple quantization steps are “compressed” into a single or few LUT entries;
+enlarge: for high-curvature regions, more LUT entries are assigned to preserve accuracy;
+unchange: for intervals that appear nearly linear, further subdivision is unnecessary.
+As a result, extra entries can be concentrated in critical intervals (e.g., [-9, 8] for SiLU) to capture rapid nonlinear variations, while intervals far from the main dynamic range (e.g., , where the function is saturated) are merged.
+Let the high-curvature region have length enlarge_length, the full input domain be , and the total number of LUT entries be table_n.
+In a single linear LUT scheme, the number of entries assigned to the high-curvature region follows:
+For example, if the SiLU function spans (hence ) and we employ a 512-entry linear LUT, this formula indicates that only entries would fall within the high-curvature region.
+By contrast, DuLUT retains the same total of 512 entries but splits them into two smaller 256-entry LUTs.
+The first LUT maps the input into a “nonlinear index,” while the second LUT stores the actual function values.
+Continuing the SiLU example, if we allocate 256 entries following the above principle, the enlarge region might occupy 4 entries, the shrink region 126 entries (with only 1 used explicitly), and thus the enlarge region ultimately gains 129 entries.
+In effect, this yields a lookup resolution equivalent to using approximately 7588 entries in a single-table design.
+A example, with 8-bit quantization, DuLUT uses two tables of 32 entries each without compromising precision (see Fig.
+8 and Algorithm 1).
 We applied DuLUT to common activation functions like softmax, GELU, and SiLU.
 By utilizing DuLUT, we achieve the same precision as larger single-table lookups while significantly reducing SRAM overhead and maintaining computational efficiency.
+| Bac | Size | Feat | Mode | mAP | NDS | mATE | mASE | mAOE | mAVE | mAAE | |
+| R50 | 512 1408 | c5 | PETR | FP32 | 31.42 | 36.11 | 84.19 | 28.42 | 60.69 | 99.08 | 23.58 |
+| PTQ | 13.79 | 25.31 | 107.94 | 31.47 | 75.22 | 83.71 | 25.45 | ||||
+| Q- PETR | FP32 | 31.49 | 37.20 | 82.52 | 27.88 | 59.91 | 91.74 | 23.45 | |||
+| PTQ | 31.34 | 37.17 | 82.61 | 27.93 | 60.00 | 91.79 | 23.45 | ||||
+| R50 | 512 1408 | p4 | PETR | FP32 | 32.60 | 37.16 | 82.63 | 27.96 | 61.06 | 95.81 | 23.91 |
+| PTQ | 12.97 | 24.75 | 108.28 | 31.76 | 79.57 | 78.90 | 27.14 | ||||
+| Q- PETR | FP32 | 32.69 | 38.03 | 80.58 | 27.89 | 59.43 | 92.69 | 22.55 | |||
+| PTQ | 32.40 | 37.72 | 81.11 | 27.92 | 60.02 | 92.76 | 22.59 | ||||
+| R101 | 512 1408 | p4 | PETR | FP32 | 34.40 | 38.62 | 80.67 | 28.03 | 57.13 | 95.74 | 24.20 |
+| PTQ | 13.53 | 23.84 | 111.04 | 31.27 | 78.94 | 92.92 | 26.14 | ||||
+| Q- PETR | FP32 | 34.72 | 39.68 | 79.40 | 27.92 | 53.90 | 92.99 | 22.59 | |||
+| PTQ | 34.41 | 39.08 | 80.98 | 28.00 | 54.41 | 92.62 | 22.70 | ||||
+| V2-99 | 320 800 | p4 | PETR | FP32 | 38.01 | 42.56 | 75.79 | 26.84 | 50.58 | 90.13 | 21.07 |
+| PTQ | 10.46 | 23.64 | 112.41 | 33.00 | 85.96 | 71.83 | 25.12 | ||||
+| Q- PETR | FP32 | 38.43 | 43.80 | 74.79 | 27.29 | 49.76 | 82.11 | 20.15 | |||
+| PTQ | 37.93 | 43.17 | 75.50 | 27.78 | 50.07 | 82.41 | 20.38 | ||||
+| V2-99 | 640 1600 | p4 | PETR | FP32 | 40.66 | 46.05 | 71.76 | 27.07 | 42.23 | 80.68 | 21.06 |
+| PTQ | 6.40 | 20.98 | 117.38 | 34.85 | 83.38 | 76.83 | 27.10 | ||||
+| Q- PETR | FP32 | 41.35 | 45.99 | 72.18 | 26.91 | 45.05 | 82.03 | 20.67 | |||
+| PTQ | 40.95 | 45.64 | 73.40 | 27.05 | 45.61 | 82.17 | 20.68 | ||||
+| V2-99 | 320 800 | p4 | StreamPETR | FP32 | 48.19 | 57.11 | 60.99 | 25.58 | 37.54 | 26.28 | 19.43 |
+| PTQ | 18.52 | 36.47 | 76.39 | 31.44 | 47.03 | 30.22 | 20.99 | ||||
+| Q- StreamPETR | FP32 | 49.13 | 57.57 | 60.77 | 26.14 | 39.05 | 24.81 | 19.15 | |||
+| PTQ | 48.21 | 56.33 | 63.00 | 26.35 | 39.17 | 24.90 | 19.19 | ||||
+| V2-99 | 640 1600 | p4 | StreamPETR | FP32 | 49.51 | 58.03 | 60.10 | 26.07 | 35.65 | 25.91 | 19.60 |
+| PTQ | 18.72 | 35.66 | 74.32 | 30.39 | 41.49 | 30.53 | 20.82 | ||||
+| Q- StreamPETR | FP32 | 50.48 | 58.61 | 58.78 | 26.16 | 37.05 | 25.69 | 18.59 | |||
+| PTQ | 49.44 | 57.94 | 60.30 | 26.55 | 37.07 | 25.87 | 18.59 | ||||
 4 Experiment
-4.1 Experimental Setup
+Detailed descriptions of benchmark, along with metric and further experimental details, are elaborated in Sup. B.
+4.1 Validation on PETR-series Methods
+We evaluate the effectiveness of our method on various PETR-series models from both FP and quantized performance perspectives, specifically considering single-frame PETR and temporal multi-frame StreamPETR models.
+Firstly, we analyze changes in floating-point performance (values in parentheses in Tab.
+2).
+In single-frame PETR models, mAP and NDS generally improve across configurations, except for a slight decrease of 0.06 in NDS when using V2-99’s P4 feature with 640×1600 resolution images.
+mAP increases range from 0.07 to 0.69, while NDS shows significant gains in all cases, ranging from 0.87 to 1.24.
+For temporal multi-frame StreamPETR models, both mAP and NDS consistently improve, with mAP gains of 0.93 and 0.94, and modest NDS increases of 0.46 and 0.58.
+Notably, NDS improvements in temporal methods are smaller than in single-frame methods, mainly due to performance degradation in mASE and mAOE, suggesting that our QDPE may not optimally capture scale and orientation information in temporal models.
+We plan to investigate this further in future work.
+Overall, QPETR shows significant improvements in most configuration metrics, demonstrating that our method surpasses the original PETR models in floating-point performance.
+Secondly, we analyze the quantized performance improvements.
+In single-frame PETR models, mAP and NDS drops are kept below 1% using our QDPE and smoothing techniques.
+In temporal multi-frame StreamPETR models, mAP and NDS drops remain within 2.5%, likely due to accumulated quantization errors during temporal fusion.
+Overall, quantized QPETR models maintain high performance with minimal drops in both settings, demonstrating the effectiveness of our quantization strategies in preserving accuracy while reducing computational and memory requirements.
+We intend to further mitigate quantization errors in temporal models through enhanced error correction or advanced quantization methods.
+4.2 Ablation Study
+Proof of Position Encoding Equivalence.
+We conducted experiments to verify whether the proposed QDPE enhances floating-point performance over the original camera-ray PE.
+As shown in Tab.
+3, QDPE provides performance improvements.
+On PETR, it slightly increases mAP by 0.07 but significantly boosts NDS and mATE by 1.09 and 1.67, respectively.
+For Stream-PETR, our method yields substantial and balanced enhancements, with increases of 0.94 in mAP, 0.46 in NDS, and 0.22 in mATE.
+| Method | mAP | NDS | mATE | |
+| PETR | Camera-ray PE | 31.42 | 36.11 | 84.19 |
+| QDPE | 31.49 | 37.20 | 82.52 | |
+| Stream -PETR | Camera-ray PE | 48.19 | 57.11 | 60.99 |
+| QDPE | 49.13 | 57.57 | 60.77 | |
+Quantization Performance Evaluation.
+We evaluate the Camera-ray PE module on the nuScenes dataset under three configurations: FP32 Baseline (full precision as an upper bound), Standard 8-bit PTQ (per-tensor 8-bit post-training quantization), and PTQ4ViT (using the PTQ4ViT [55] method to boost accuracy).
+As shown in Table 4, retaining the Softmax input in full precision (“No”) yields higher mAP and NDS than when it is quantized (“Yes”), underscoring the importance of careful Softmax treatment.
+| Method | Quant. Softmax Input | (PTQ) INT8 | (PTQ4ViT) INT8 | ||
+| mAP | NDS | mAP | NDS | ||
+| Camera- ray PE | FP32 | 31.42 | 36.11 | 31.42 | 36.11 |
+| No | 24.90 | 32.10 | 27.10 | 33.60 | |
+| Yes | 18.80 | 27.50 | 19.20 | 28.40 | |
+Compare with QAT.
+Although QDPE requires retraining, its cost is comparable to that of QAT.
+We implement a distillation-based QAT (inspired by QD-BEV [57]) on the original Camera-ray PE.
+As shown in Table 5, even after 24 or 36 epochs, QAT yields lower mAP and NDS than QDPE with standard PTQ.
+This confirms that our amplitude-aware design not only maintains high floating-point performance but also achieves superior quantized accuracy.
+| Model | 12 epochs | 24 epochs | 36 epochs | ||||
+| mAP | NDS | mAP | NDS | mAP | NDS | ||
+| Camera-ray PE | QAT (Distill) | 28.9 | 35.2 | 30.3 | 35.8 | 30.3 | 35.8 |
+| QDPE | PTQ | 31.34 | 37.17 | 31.34 | 37.17 | 31.34 | 37.17 |
+Effect of Anchor Embedding Quantity.
+The QDPE uses three anchor embeddings per axis, obtained through linear interpolation.
+Experiments (Tab.
+6) demonstrate that setting the number of anchor embeddings to 3 achieves the highest NDS and mAP scores.
+Adjusting this number either up or down results in lower performance, confirming that 3 is the optimal choice.
+| Quantity of Anchor Embedding | NDS | mAP | ||
+| x-axis | y-axis | z-axis | ||
+| 2 | 2 | 2 | 36.66 | 31.29 |
+| 3 | 3 | 3 | 37.20 | 31.49 |
+| 4 | 4 | 4 | 36.92 | 31.09 |
+| 5 | 5 | 5 | 36.83 | 31.19 |
+Quantization Performance of Different Position Encodings.
+To experimentally demonstrate the superior quantization performance of our proposed QDPE, we focus solely on quantizing the positional encoding, keeping all other modules in floating-point computation.
+Detailed results are shown in Tab.
+7.
+The original camera-ray configuration loses up to 11.97% in mAP and 5.04% in NDS, whereas our QDPE experiences minimal losses of only 1.42% in mAP and 1.15% in NDS.
+Fig.
+9 further supports this finding; compared to the distribution in Fig.
+4, the distribution of our QDPE aligns more closely with that of image features, retaining sufficient useful information.
+| Method | NDS | mAP | mATE | |
+| PETR | Camera-ray PE | 34.29 | 27.66 | 87.17 |
+| QDPE | 37.18 | 31.40 | 82.59 | |
+| Stream -PETR | Camera-ray PE | 53.74 | 40.23 | 69.39 |
+| QDPE | 56.81 | 47.65 | 61.53 | |
+Impact of Different Scaled Dot Product Quantization Strategy.
+To validate our novel scaled dot-product quantization strategy—which searches for the optimal scaling factor by minimizing softmax output error—we focus solely on quantizing the softmax inputs while keeping all other modules in floating-point computation.
+As shown in Tab.
+8, the original quantization strategy results in significant losses of 40% in NDS and 50% in mAP.
+In contrast, our ”quant after stabilization” approach greatly improves performance.
+An ablation study on the maximum candidate truncation range reveals that setting yields optimal quantization performance with nearly no loss.
+Performance deteriorates when due to increased truncation of feature information, while values of exceeding 20 offer no additional benefits.
+Therefore, setting is sufficient to achieve the best performance.
+Additionally, in large language models, the attention inputs can reach extremely large values (see Fig.
+10), and we have validated the effectiveness of our method in such scenarios as well (see Tab.
+9).
+| Method | NDS | mAP | mATE | |
+| quant before ns | - | 25.31 | 13.79 | 107.94 |
+| quant after ns | N = 10 | 3.45 | 1.23 | 150.34 |
+| N = 50 | 33.86 | 28.77 | 87.12 | |
+| N = 10 | 34.65 | 29.33 | 85.01 | |
+| N = 20 | 36.10 | 31.42 | 84.19 | |
+| N = 30 | 36.10 | 31.42 | 84.19 | |
+| N = 40 | 36.10 | 31.42 | 84.19 | |
+| Model Name | qwen2.5-7b-instruct | deepseek-r1-distill-qwen-7b | ||
+| wikitext2 | gsm8k | wikitext2 | gsm8k | |
+| bfp16 | 7.46 | 80.21 | 25.04 | 85.97 |
+| quant before ns | 10000+ | 0.3 | 10000+ | 0.1 |
+| quant after ns(20) | 7.48 | 80.24 | 25.09 | 86.03 |
+Superior Performance of DuLUT for Non-linear Functions.
+To validate the quantization advantages of our proposed DuLUT for nonlinear functions, we use ”quant after stabilization (N=20)” from Tab.
+8 as a baseline and evaluate the performance with different nonlinear function quantization methods applied on top of it.
+The specific results are shown in Tab.
+10.
+We consider the carefully designed approximation methods I-Bert and I-Vit for different nonlinear functions.
+Due to the approximation errors introduced by these methods, many points are quantized away.
+Additionally, we compare with the LUT-based table lookup method and find that 256 entries are required for lossless quantization, while 128 entries lead to severe performance losses of 0.54 NDS and 0.37 mAP.
+In contrast, our newly proposed DuLUT with 128 entries achieves lossless quantization.
+Even when the number of entries is further reduced to 64, the quantization only results in a negligible loss of 0.08% NDS and 0.02% mAP, which can be considered negligible.
+This further demonstrates the superior quantization performance of our proposed DuLUT.
+| Method | NDS | mAP | mATE | |
+| Quant after stabilization (N=20) | 36.10 | 31.42 | 84.19 | |
+| I-Bert | 34.87 | 29.34 | 88.41 | |
+| I-Vit | 35.03 | 28.77 | 87.32 | |
+| LUT | 256 entries | 36.10 | 31.42 | 84.19 |
+| 128 entries | 35.56 | 31.05 | 85.61 | |
+| DuLUT | (16,16) entries | 28.12 | 17.36 | 96.99 |
+| (16,32) entries | 34.14 | 27.29 | 90.33 | |
+| (32,32) entries | 36.07 | 31.36 | 84.20 | |
+| (64,64) entries | 36.10 | 31.42 | 84.19 | |
+Practical Hardware Resource Savings.
+Tab.
+11 shows Q-PETR runs at 13.3 FPS (87% faster) and 1.9 GB memory (60% less) vs.
+PETR’s 7.1 FPS/4.8 GB, demonstrating significant speedup and resource efficiency.
+| Method | Mode | FPS | CUDA memory (G) |
+| PETR | fp32 | 7.1 | 4.8 |
+| Q-PETR | INT8 | 13.3 | 1.9 |
+5 Conclusion
+In this paper, we address the significant performance drops of PETR models during quantization by identifying two main issues: the imbalance between positional encoding and image feature magnitudes, and uneven scalar dot-products in cross-attention.
+To resolve these, we introduce Q-PETR, a quantization-friendly positional encoding transformation that redesigns positional encoding and improves scalar dot-product quantization without sacrificing the original floating-point performance.
+We also propose DuLUT, a dual-table lookup mechanism for efficiently quantizing nonlinear functions, further enhancing deployment suitability on edge AI chips.
+Our experiments show that Q-PETR limits mAP and NDS drops to below 1% under standard 8-bit post-training quantization and even surpasses the original PETR in floating-point precision.
+Extensive tests across various PETR models demonstrate the method’s strong generalization and deployment suitability.
+Supplementary
+Appendix A Preliminaries
+PETR
+enhances 2D image features with 3D position-aware properties using camera-ray positional encoding (PE), enabling refined query updates for 3D bounding box prediction.
+Specifically, surround-view images pass through a backbone to generate 2D features , while camera-ray PE is computed using camera intrinsics and extrinsics.
+The learnable query embeddings serve as the initial queries for the decoder.
+Here, serves as the values , and adding to element-wise forms the 3D position-aware keys .
+The decoder updates the queries using these key-value pairs through self-attention, cross-attention, and feed-forward network (FFN) modules.
+The updated query vectors are passed through an MLP to predict 3D bounding box categories and attributes, repeating for cycles.
+The entire PETR process is summarized in Algorithm 2.
+Appendix B Experimental Setup
 Benchmark.
@@
 When quantizing the scaled dot-product in cross-attention, we define a candidate set of 20 scaling factors.
-4.2 Ablation Study
-Proof of Position Encoding Equivalence We conducted experiments to verify whether the proposed QD-aware lidar-ray PE enhances floating-point performance over the original camera-ray PE.
-As shown in Tab.
-2, QD-aware lidar-ray PE provides performance improvements.
-On PETR, it slightly increases mAP by 0.07 but significantly boosts NDS and mATE by 1.09 and 1.67, respectively.
-For Stream-PETR, our method yields substantial and balanced enhancements, with increases of 0.94 in mAP, 0.46 in NDS, and 0.22 in mATE.
-| Method | mAP | NDS | mATE | |
-| PETR | Camera-ray PE | 31.42(0.00) | 36.11(0.00) | 84.19(0.00) |
-| QD-aware lidar-ray PE | 31.49(0.07) | 37.20(1.09) | 82.52(1.67) | |
-| Stream -PETR | Camera-ray PE | 48.19(0.00) | 57.11(0.00) | 60.99(0.00) |
-| QD-aware lidar-ray PE | 49.13(0.94) | 57.57(0.46) | 60.77(0.22) | |
-Effect of Anchor Embedding Quantity.
-The QD-aware lidar-ray PE uses three anchor embeddings per axis, obtained through linear interpolation.
-Experiments (Tab.
-3) demonstrate that setting the number of anchor embeddings to 3 achieves the highest NDS and mAP scores.
-Adjusting this number either up or down results in lower performance, confirming that 3 is the optimal choice.
-| Quantity of Anchor Embedding | NDS | mAP | ||
-| x-axis | y-axis | z-axis | ||
-| 2 | 2 | 2 | 36.66 | 31.29 |
-| 3 | 3 | 3 | 37.20 | 31.49 |
-| 4 | 4 | 4 | 36.92 | 31.09 |
-| 5 | 5 | 5 | 36.83 | 31.19 |
-Quantization Performance of Different Position Encodings To experimentally demonstrate the superior quantization performance of our proposed QD-aware lidar-ray PE, we focus solely on quantizing the positional encoding, keeping all other modules in floating-point computation.
-Detailed results are shown in Tab.4.
-The original camera-ray configuration loses up to 11.97% in mAP and 5.04% in NDS, whereas our QD-aware lidar-ray PE experiences minimal losses of only 1.42% in mAP and 1.15% in NDS.
-Figure8 further supports this finding; compared to the distribution in Fig.
-3, the distribution of our QD-aware lidar-ray PE aligns more closely with that of image features, retaining sufficient useful information.
-| Method | NDS | mAP | mATE | |
-| PETR | Camera-ray PE | 34.29 | 27.66 | 87.17 |
-| QD-aware lidar-ray PE | 36.77 | 31.04 | 82.92 | |
-| Stream -PETR | Camera-ray PE | 53.74 | 40.23 | 69.39 |
-| QD-aware lidar-ray PE | 56.81 | 47.65 | 61.53 | |
-Impact of Different Scaled Dot Product Quantization Strategy.
-To validate our novel scaled dot-product quantization strategy—which searches for the optimal scaling factor by minimizing softmax output error—we focus solely on quantizing the softmax inputs while keeping all other modules in floating-point computation.
-As shown in Tab.
-5, the original quantization strategy results in significant losses of 40% in NDS and 50% in mAP.
-In contrast, our ”quant after stabilization” approach greatly improves performance.
-An ablation study on the maximum candidate truncation range reveals that setting yields optimal quantization performance with nearly no loss.
-Performance deteriorates when due to increased truncation of feature information, while values of exceeding 20 offer no additional benefits.
-Therefore, setting is sufficient to achieve the best performance.
-| Method | NDS | mAP | mATE | |
-| quant before stabilization | - | 25.31 | 13.79 | 107.94 |
-| quant after stabilization | N = 10 | 3.45 | 1.23 | 150.34 |
-| N = 50 | 33.86 | 28.77 | 87.12 | |
-| N = 10 | 34.65 | 29.33 | 85.01 | |
-| N = 20 | 36.10 | 31.42 | 84.19 | |
-| N = 30 | 36.10 | 31.42 | 84.19 | |
-| N = 40 | 36.10 | 31.42 | 84.19 | |
-| Backbone | Input Size | Feature | Mode | mAP | NDS | mATE | mASE | mAOE | mAVE | mAAE | |
-| R50 | 5121408 | c5 | PETR | FP32 | 31.42(0.00) | 36.11(0.00) | 84.19(0.00) | 28.42(0.00) | 60.69(0.00) | 99.08(0.00) | 23.58(0.00) |
-| PTQ | 13.79[56.11] | 25.31[29.91] | 107.94[28.21] | 31.47[10.73] | 75.22[23.94] | 83.71[15.51] | 25.45[7.93] | ||||
-| Q-PETR | FP32 | 31.49(0.07) | 37.20(1.09) | 82.52(1.67) | 27.88(0.54) | 59.91(0.78) | 91.74(7.34) | 23.45(0.13) | |||
-| PTQ | 31.34[0.47] | 37.17[0.82] | 82.61[0.65] | 27.93[0.17] | 60.00[0.15] | 91.79[0.00] | 23.45[0.00] | ||||
-| R50 | 5121408 | p4 | PETR | FP32 | 32.60(0.00) | 37.16(0.00) | 82.63(0.00) | 27.96(0.00) | 61.06(0.00) | 95.81(0.00) | 23.91(0.00) |
-| PTQ | 12.97[60.21] | 24.75[33.39] | 108.28[31.04] | 31.76[13.59] | 79.57[30.31] | 78.90[17.65] | 27.14[13.51] | ||||
-| Q-PETR | FP32 | 32.69(0.09) | 38.03(0.87) | 80.58(2.05) | 27.89(0.07) | 59.43(0.63) | 92.69(3.12) | 22.55(1.36) | |||
-| PTQ | 32.40[0.88] | 37.72[0.82] | 81.11[0.65] | 27.92[0.10] | 60.02[0.97] | 92.76[0.00] | 22.59[0.18] | ||||
-| R101 | 5121408 | p4 | PETR | FP32 | 34.40(0.00) | 38.62(0.00) | 80.67(0.00) | 28.03(0.00) | 57.13(0.00) | 95.74(0.00) | 24.20(0.00) |
-| PTQ | 13.53[60.67] | 23.84[38.27] | 111.04[37.65] | 31.27[11.56] | 78.94[38.18] | 92.92[2.95] | 26.14[8.02] | ||||
-| Q-PETR | FP32 | 34.72(0.32) | 39.68(1.06) | 79.40(1.27) | 27.92(0.11) | 53.90(3.23) | 92.99(2.75) | 22.59(1.61) | |||
-| PTQ | 34.41[0.89] | 39.08[1.51] | 80.98[1.98] | 28.00[0.28] | 54.41[0.94] | 92.62[0.39] | 22.70[0.47] | ||||
-| V2-99 | 320800 | p4 | PETR | FP32 | 38.01(0.00) | 42.56(0.00) | 75.79(0.00) | 26.84(0.00) | 50.58(0.00) | 90.13(0.00) | 21.07(0.00) |
-| PTQ | 10.46[72.48] | 23.64[44.45] | 112.41[36.21] | 33.00[22.95] | 85.96[69.95] | 71.83[20.30] | 25.12[19.22] | ||||
-| Q-PETR | FP32 | 38.43(0.42) | 43.80(1.24) | 74.79(1.00) | 27.29(0.45) | 49.76(0.82) | 82.11(8.02) | 20.15(0.92) | |||
-| PTQ | 37.93[1.30] | 43.17[1.44] | 75.50[0.94] | 27.78[1.79] | 50.07[0.62] | 82.41[0.36] | 20.38[1.14] | ||||
-| V2-99 | 6401600 | p4 | PETR | FP32 | 40.66(0.00) | 46.05(0.00) | 71.76(0.00) | 27.07(0.00) | 42.23(0.00) | 80.68(0.00) | 21.06(0.00) |
-| PTQ | 6.40[84.63] | 20.98[54.55] | 117.38[33.22] | 34.85[25.92] | 83.38[97.44] | 76.83[4.79] | 27.10[28.67] | ||||
-| Q-PETR | FP32 | 41.35(0.69) | 45.99(0.06) | 72.18(0.42) | 26.91(0.15) | 45.05(2.82) | 82.03(2.35) | 20.67(0.39) | |||
-| PTQ | 40.95[0.96] | 45.64[1.09] | 73.40[1.69] | 27.05[0.52] | 45.61[1.24] | 82.17[0.17] | 20.68[0.00] | ||||
-| V2-99 | 320800 | p4 | StreamPETR | FP32 | 48.19(0.00) | 57.11(0.00) | 60.99(0.00) | 25.58(0.00) | 37.54(0.00) | 26.28(0.00) | 19.43(0.00) |
-| PTQ | 18.52(61.19) | 36.47(36.11) | 76.39(25.25) | 31.44 (22.90) | 47.03 (25.27) | 30.22 14.99 | 20.99 (8.03) | ||||
-| Q-StreamPETR | FP32 | 49.13(0.94) | 57.57(0.46) | 60.77(0.22) | 26.14(0.56) | 39.05(1.49) | 24.81(1.47) | 19.15(0.28) | |||
-| PTQ | 48.21 (1.87) | 56.33 (2.15) | 63.00 (3.66) | 26.35 (0.80) | 39.17 (0.31) | 24.90 (0.36) | 19.19 (0.21) | ||||
-| V2-99 | 6401600 | p4 | StreamPETR | FP32 | 49.51(0.00) | 58.03(0.00) | 60.10(0.00) | 26.07(0.00) | 35.65(0.00) | 25.91(0.00) | 19.60(0.00) |
-| PTQ | 18.72(62.19) | 35.66(38.54) | 74.32(23.66) | 30.39 (16.76) | 41.49 (16.38) | 30.53 17.83 | 20.82 (6.22) | ||||
-| Q-StreamPETR | FP32 | 50.48(0.93) | 58.61(0.58) | 58.78(0.32) | 26.16(0.09) | 37.05(1.40) | 25.69(0.22) | 18.59(1.01) | |||
-| PTQ | 49.44 (2.06) | 57.94 (1.24) | 60.30 (2.52) | 26.55 (1.49) | 37.07 (0.00) | 25.87 (0.31) | 18.59 (0.00) | ||||
-Superior Performance of DuLUT for Non-linear Functions.
-To validate the quantization advantages of our proposed DuLUT for nonlinear functions, we use ”quant after stabilization (N=20)” from Tab.
-5 as a baseline and evaluate the performance with different nonlinear function quantization methods applied on top of it.
-The specific results are shown in Tab.
-7.
-We consider the carefully designed approximation methods I-Bert and I-Vit for different nonlinear functions.
-Due to the approximation errors introduced by these methods, many points are quantized away.
-Additionally, we compare with the LUT-based table lookup method and find that 256 entries are required for lossless quantization, while 128 entries lead to severe performance losses of 0.54 NDS and 0.37 mAP.
-In contrast, our newly proposed DuLUT with 128 entries achieves lossless quantization.
-Even when the number of entries is further reduced to 64, the quantization only results in a negligible loss of 0.08% NDS and 0.02% mAP, which can be considered negligible.
-This further demonstrates the superior quantization performance of our proposed DuLUT.
-| Method | NDS | mAP | mATE | |
-| Quant after stabilization (N=20) | 36.10 | 31.42 | 84.19 | |
-| I-Bert | 34.87 | 29.34 | 88.41 | |
-| I-Vit | 35.03 | 28.77 | 87.32 | |
-| LUT | 256 entries | 36.10 | 31.42 | 84.19 |
-| 128 entries | 35.56 | 31.05 | 85.61 | |
-| DuLUT | (16,16) entries | 28.12 | 17.36 | 96.99 |
-| (16,32) entries | 34.14 | 27.29 | 90.33 | |
-| (32,32) entries | 36.07 | 31.36 | 84.20 | |
-| (64,64) entries | 36.10 | 31.42 | 84.19 | |
-4.3 Validation on PETR-series Methods
-We evaluate the effectiveness of our proposed method on various PETR-series models from both floating-point and quantized performance perspectives, specifically considering single-frame PETR and temporal multi-frame StreamPETR models.
-First, we analyze changes in floating-point performance (values in parentheses in Tab.
-6).
-In single-frame PETR models, mAP and NDS generally improve across configurations, except for a slight decrease of 0.06 in NDS when using V2-99’s P4 feature with 640×1600 resolution images.
-mAP increases range from 0.07 to 0.69, while NDS shows significant gains in all cases, ranging from 0.87 to 1.24.
-For temporal multi-frame StreamPETR models, both mAP and NDS consistently improve, with mAP gains of 0.93 and 0.94, and modest NDS increases of 0.46 and 0.58.
-Notably, NDS improvements in temporal methods are smaller than in single-frame methods, mainly due to performance degradation in mASE and mAOE, suggesting that our QD-aware lidar-ray PE may not optimally capture scale and orientation information in temporal models.
-We plan to investigate this further in future work.
-Overall, QPETR shows significant improvements in most configuration metrics, demonstrating that our method surpasses the original PETR models in floating-point performance.
-Next, we analyze the quantized performance improvements (values in square brackets in Table 6).
-In single-frame PETR models, mAP and NDS drops are kept below 1% using our QD-aware lidar-ray PE and smoothing techniques.
-In temporal multi-frame StreamPETR models, mAP and NDS drops remain within 2.5%, likely due to accumulated quantization errors during temporal fusion.
-Overall, quantized QPETR models maintain high performance with minimal drops in both settings, demonstrating the effectiveness of our quantization strategies in preserving accuracy while reducing computational and memory requirements.
-We intend to further mitigate quantization errors in temporal models through enhanced error correction or advanced quantization methods.
-5 Limitations
-Although our method incurs almost no quantization accuracy loss, users need to replace the camera-ray in the original PETR series with our proposed QD-aware lidar-ray.
+Appendix C Theoretical Analysis of Magnitude Bounds in Position Encodings
+C.1 Normalization Framework and Input Conditioning
+To establish a unified analytical framework, we first formalize the spatial normalization process for various ray-based position encodings.
+Let denote the 3D coordinates within the perception range meters and meters.
+The normalized coordinates are computed as:
+Noting that is clamped to within the range , the distribution ranges of the normalized sampled points in positional encodings are characterized as follows:
+For the sampled point of Camera-Ray PE, denoted as , the distribution spans the unit cube, i.e., .
+For the sampled points of LiDAR-Ray PE and QDPE, denoted as and respectively, the distributions are constrained to .
+Here, the value is derived from the ratio , where corresponds to the fixed depth setting in the encoding process.
+This distinction highlights the inherent differences in spatial coverage and normalization strategies employed by these positional encodings.
+C.2 Magnitude Propagation Analysis
+C.2.1 Camera-Ray Position Encoding
+As illustrated in Fig. 6 (a), the encoding pipeline consists of two critical stages:
+Stage 1: Inverse Sigmoid Transformation
+Empirical analysis reveals a maximum magnitude .
+Stage 2: MLP Projection (Through Two Fully-Connected Layers)
+where denotes the ReLU activation function. Let be the maximum weight magnitude. We derive the upper bound:
+where and denote the input tensor channels for and , respectively.
+C.2.2 LiDAR-Ray Position Encoding
+Unlike Camera-Ray PE, the encoding process of LiDAR-Ray PE introduces sinusoidal modulation between the inverse sigmoid transformation and MLP projection, as shown in Fig.
+6 (b).
+The magnitude propagation for LiDAR-Ray PE is as follows:
+Stage 1: Inverse Sigmoid Transformation
+Empirical analysis reveals a maximum magnitude .
+Stage 2: Spectral Embedding
+where denotes concatenation. This ensures:
+Stage 3: MLP Projection (Following setting in Camera-Ray PE)
+C.2.3 Ours QD-PE
+The proposed encoding introduces anchor-based constraints, as depicted in Fig. 6 (c):
+Stage 1: Anchor Interpolation (For Each Axis )
+where denotes learnable anchor embeddings. Via Theorem C.1, the magnitude is constrain to:
+Stage 2: MLP Projection
+C.3 Comparative Magnitude Analysis
+The derived bounds reveal fundamental differences in magnitude scaling:
+This analysis demonstrates that QD-PE requires less quantization range than Camera-Ray PE.
+C.4 Theoretical Guarantee of Magnitude Constraints
+Theorem C.1 (Anchor Embedding Magnitude Bound).
+Let be adjacent anchor embeddings with . For any point , its interpolated embedding satisfies:
+Proof.
+Let . The interpolated embedding becomes:
+For any component :
+Thus, holds for all dimensions. ∎
+Through the application of regularization (e.g., L2 constraint) on the anchor embeddings during training, the magnitude of can be explicitly controlled.
+Empirically, we find that this value converges to approximately 0.8 in our experiments.
+Appendix D More Ablation Study
+D.1 Local Similarity of Position Encoding Features
+Fig.
+11 shows that QD-PE significantly outperforms 3D point PE and cameraray PE in local similarity of position encoding.
+Its similarity distribution appears more compact and concentrated, validating the method’s superiority in local spatial information modeling and its capability to precisely capture neighborhood spatial relationships around target pixels.
+Appendix E Limitations
+Although our method incurs almost no quantization accuracy loss, users need to replace the camera-ray in the original PETR series with our proposed QDPE.
 The only drawback is that this requires retraining.
 However, from the perspective of quantization deployment, this retraining is beneficial, and the floating-point precision can even be improved.
-6 Conclusion
-In this paper, we address the significant performance drops of PETR models during quantization by identifying two main issues: the imbalance between positional encoding and image feature magnitudes, and uneven scalar dot-products in cross-attention.
-To resolve these, we introduce Q-PETR, a quantization-friendly positional encoding transformation that redesigns positional encoding and improves scalar dot-product quantization without sacrificing the original floating-point performance.
-Our experiments show that Q-PETR limits mAP and NDS drops to below 1% under standard 8-bit post-training quantization and even surpasses the original PETR in floating-point precision.
-Extensive tests across various PETR models demonstrate the method’s strong generalization and deployment suitability.
-Future work will explore quantization error propagation in multi-frame temporal models and develop advanced quantization strategies to further reduce precision loss in temporal settings.